OpenStudy (girlgoyleh.):

What is the slope of the line that passes through the points (14, 5) and (20, –4)?

OpenStudy (anonymous):

Use the slope formula :

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 14}}\quad ,&{\color{blue}{ 5}})\quad &({\color{red}{ 20}}\quad ,&{\color{blue}{ -4}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (luigi0210):

Use the slope formula first: \[\LARGE m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] Then point slope: \[\LARGE y-y_{1}=m(x-x_{1}\]

OpenStudy (luigi0210):

Lost a bracket there o:

OpenStudy (anonymous):

M=(20-14)/(-4-5)

OpenStudy (girlgoyleh.):

So it would be -3/2

OpenStudy (anonymous):

-2/3? because -6/9

OpenStudy (girlgoyleh.):

Thanks for your guys' help! It was -3/2

OpenStudy (anonymous):

Nevermind me then D: